Hi! I need help solving a word problem, a step by step demo would be appreciated and helpful!
Here's the problem:
Liz and Troy collect baseball cards. Liz has 4 less than twice the number of cards Troy has. Together they have 149 cards. How many does each person have?
This has me stuck! Thanks in advance :)

Found 2 solutions by josgarithmetic, greenestamps:
Answer by josgarithmetic(39800)

LIZ           2y-4

TROY           y

total         149

Answer by greenestamps(13334)



Step 1: Using the given information, define variable(s) and expression(s) to represent the numbers in the problem.

Let x = number Troy has
Then 2x-4 = number Liz has (4 less than twice as many as Troy)

Step 2: Use those variables and expressions to write an equation based on the given information.

(x) + (2x-4) = 149 (the total number of cards the two of them have is 149)

Step 3: Solve the equation using basic algebra to find x

Step 4: Use the value of x you found to determine the number of cards each one has.

Step 5: Verify that your answer satisfies all the conditions of the problem.

I leave it to you to do steps 3 to 5....